OpenStudy (anonymous):

solve -15(4+3x)-8x=2b-6(x-9). please show how you got this

sam (.sam.):

\[\begin{array}{l} -8 x-15 (3 x+4)=2 b-6 (x-9) \\ \text{Expand out terms on both sides} \\ -53 x-60=2 b-6 x+54 \\ \text{Add }(6 x+60)\text{ to both sides} \\ -47 x=2 b+114 \\ \text{Divide both sides by }-47 \\ x=\frac{(-114-2 b)}{47} \\\end{array}\]

OpenStudy (anonymous):

woah how did you get that your good